# Adaptive Collaboration Strategy for LLMs in Medical Decision Making

### v1

@@ Abstract
Foundation models have become invaluable in advancing the medical field. Despite their promise, the strategic deployment of LLMs for effective utility in complex medical tasks remains an open question. Our novel framework, Medical Decision-making Agents (MDAgents) aims to address this gap by automatically assigning the effective collaboration structure for LLMs. Assigned solo or group collaboration structure is tailored to the complexity of the medical task at hand, emulating real-world medical decision making processes. We evaluate our framework and baseline methods with state-of-the-art LLMs across a suite of challenging medical benchmarks: MedQA, MedMCQA, PubMedQA, DDXPlus, PMC-VQA, Path-VQA, and MedVidQA, achieving the best performance in 5 out of 7 benchmarks that require an understanding of multi-modal medical reasoning. Ablation studies reveal that MDAgents excels in adapting the number of collaborating agents to optimize efficiency and accuracy, showcasing its robustness in diverse scenarios. We also explore the dynamics of group consensus, offering insights into how collaborative agents could behave in complex clinical team dynamics. Our code can be found at https://github.com/mitmedialab/MDAgents
@@ 1 Introduction
Medical Decision-Making (MDM) is a multifaceted and intricate process, where clinicians navigate vast and diverse sources of information to arrive at precise conclusions under complexity [88]. Recently, Large Language Models (LLMs) have shown potential in transforming MDM [64, 16, 57] by digesting vast amounts of medical literature [67] and clinical information [1], thereby supporting probabilistic [85] and causal [32] reasoning processes crucial to medical practice. The severe implications of inaccuracies in healthcare are misdiagnoses [45] and inappropriate treatments [79, 9], which demand a uniquely careful and precise approach. Additionally, MDM involves the interpretation of complex and multi-modal data, such as imaging [5], electronic health records (EHR) [56], signals [39, 20], genetic information [55], and the rapid integration of new medical research into clinical practice [70, 61], highlighting the unique challenges in this field.
Figure 1: MDAgents: Our medical agents decision making framework. Given a medical query consisting of patient information and medical contexts, the framework performs 1) medical complexity check, 2) interaction type selection, 3) prompting/collaboration, and 4) decision making. Across multiple benchmarks that include multi-modal medical query, our framework performed the best performance in 5 out of 7 benchmarks.
@@
While decision-making tools including multi-agent systems [80, 7] have shown promise in various general domains [26, 35, 37, 58, 27, 54], their performance in healthcare has been limited. This limitation arises due to their generalist design, which lacks the specialized, in-depth medical knowledge and processes essential for accurate MDM [44]. In contrast, human clinicians apply an adaptive, collaborative, and tiered approach to MDM – considering the current and past history of the patient, available evidence from medical literature, and the clinicians’ clinical expertise and experience [14]. For instance, patients are triaged in urgent care units by the severity and complexity of their medical conditions [81, 8, 19]. Examples of low-complexity cases include patients with single uncomplicated acute cases or stable chronic conditions that could be sent to their primary care physicians (PCP) [78]. In contrast, examples of high-complexity cases can be patients with complicated injuries or chronic conditions, including side effects of treatment or with superimposed diseases. These cases often require consultation with specialty physicians [53, 21]. Inspired by this real-world MDM process, we propose Medical Decision-making Agents (MDAgents), an adaptive medical decision-making framework that leverages LLMs to emulate the hierarchical diagnosis procedures ranging from individual clinicians to collaborative clinician teams (Figure 1)
MDAgents unfolds in four stages: 1) Medical Complexity Check - The system evaluates the intricacies of the medical query, stratifying the problem into low, moderate, or high categories based on the query’s medical context. 2) Expert Recruitment - Depending on the identified complexity level, the framework activates the appropriate diagnostic method. A simple solo approach is employed for issues with low complexity, while a Multidisciplinary Team (MDT) or Integrated Care Team (ICT) is convened for more scenarios with moderate or high complexities. 3) Inference Process - For solo queries, the framework uses a set of prompting techniques (e.g. Chain-of-Thought (CoT), Self-Consistency (SC) and Medprompt) to provide answers. In cases of an MDT, multiple LLM agents with specialized medical expertise are brought together to form a consensus through collaborative discussion. For the most complex cases, an ICT synthesizes information from diverse domains to produce a comprehensive report, culminating in the final decision.
MDAgents is the first adaptive decision-making framework for LLMs that mirrors real-world MDM processes, allowing for dynamic collaboration among AI agents based on task complexity. We conduct experiments on seven medical question-answering datasets, including 1) text-only: MedQA [29], MedMCQA [52], PubMedQA [30], DDxPlus [65] 2) image + text: PMC-VQA [86], Path-VQA [25], and 3) video + text: MedVidQA [23]. Our framework demonstrates superior performance in accuracy over previous solo and group methods on 5 out of 7 medical benchmarks. Further analysis reveals that our framework can provide effective trade-off between the performance and the number of API calls by varying number of agents. Moreover, rigorous testing under various temperatures has demonstrated better robustness of our system than solo and group methods. Finally, we provide ablation studies evidencing that our framework finds the appropriate complexity level for each MDM instance.
@@
Table 1: Comparison between our MDAgents framework and other multi-agent collaboration methods. Among these works, MDAgents is the only one to perform all key dimensions of LLM decision making.
@@
Method MDAgents (Ours) Single Voting [75] Delphi [64] Debate [12] ReConcile [6]
@@
2 Related Works
2.1 LLMs in Medical Decision Making
Large language models (LLMs) have been used in the medical field for a range of medical tasks and clinical applications [68, 87]. They can answer questions from medical exams [34, 42], biomedical research [30], and clinical diagnosis [59, 43]. Some medical LLMs are also evaluated on generative tasks such as generating medical reports [71], describing medical images [73], and performing diagnostic dialogue with patients [69]. To advance the capabilities of medical LLMs, two main approaches have been explored: (1) training with domain-specific data [22], and (2) applying inference-time techniques such as prompt engineering [60] and Retrieval Augmented Generation (RAG) [83]. While initial research concentrated on pre-training and fine-tuning with medical knowledge, the rise of large general-purpose LLMs has enabled training-free methods where models leverage their latent medical knowledge. For example, GPT-4 [50], without any specialized prompt crafting, surpasses the passing score on USMLE by over 20 points and outperforms fine-tuned models including Med-PaLM [48, 47]. The promise of general-purpose models has thus inspired various techniques such as Medprompt and ensemble refinement to improve LLM reasoning [60], as well as RAG tools that use external resources to improve the factuality and completeness of LLM responses [83, 31]. Our approach leverages these techniques and the capabilities of general-purpose models, while acknowledging that a solitary LLM may not fully encapsulate the collaborative and multidisciplinary nature of real-world MDM. We thus emphasize joining multiple expert LLMs for effective collaboration in order to solve medical tasks with greater accuracy.
2.2 Multi-Agent Collaboration
An array of studies have explored effective collaboration frameworks between multiple LLM agents [38, 80] to enhance the capabilities of individual LLMs [72]. A common framework is role-playing, where each agent adopts a specific role (e.g. an Assistant Agent or a User Agent) and break down a task into sub-steps to solve it collaboratively [38, 80]. While role-playing focuses on collaboration and multi-step problem-solving [82], another framework called “multi-agent debate" prompts each agent to solve the task independently [12]. Then, they reason through other agents’ answers to converge on a shared response in order to improve factuality and performance in mathematical and reasoning tasks [12, 40]. Similar frameworks include voting [75], Delphi consensus [64], ReConcile group discussions, and negotiating [17]. Table 1 compares existing setups across key dimensions in multi-agent interaction. Although these frameworks have shown improvement in the respective tasks, they rely on a pre-determined number of agents and interaction settings. When applied to settings with a wider variety of tasks, such as medical question-answering [87], this static architecture may lead to suboptimal multi-agent configuration for certain tasks and result in limited performance [41]. Additionally, it could be computationally inefficient and expensive to employ multiple agents for simpler tasks without noticeable performance gains [12]. Given that different models and frameworks could generalize better to different tasks [84], we propose a framework that dynamically assigns the optimal collaboration strategy at inference time based on the complexity of the query. We target our strategy at the relatively untapped field of MDM, which requires a team effort and is believed to benefit from multi-agent collaboration [64].
3 Methods
Algorithm 1 Adaptive Medical Decision Making Framework
1:Problem $Q$
2:$Complexity\leftarrow\textsc{ComplexityCheck}(Q)$triangleright Determine the complexity of the medical input
3:if $Complexity=low$ then
4:    $Agent\leftarrow\textsc{Recruit}(Q,Complexity)$triangleright Recruit a single agent
5:    $ans\leftarrow Agent(Q)$
6:
7:else if $Complexity=moderate$ then
8:    $MDT\leftarrow\textsc{Recruit}(Q,Complexity)$triangleright Recruit a Multi-disciplinary Team
9:    $Agent\leftarrow\textsc{Recruit}(Q,Complexity,MDT)$
10:    $r\leftarrow 0$
11:    $Consensus\leftarrow\text{False}$
12:    $Interaction\leftarrow[]$
13:    while $r\leq R$ and not $Consensus$ do
14:         $Consensus,Log\leftarrow\textsc{CollaborativeDiscussion}(Q,MDT)$triangleright Iterative discussions
15:         if not $Consensus$ then
16:             for all $Agent\in MDT$ do
17:                 $Feedback\leftarrow Moderator(Interaction,Agent)$triangleright Moderator reviews and provides feedback
18:                 $Agent.\textsc{Update}(Feedback)$triangleright Update agent with the feedback
19:             end for
20:             $Interaction\leftarrow Interaction+[Log]+[Feedback]$
21:         end if
22:         $r\leftarrow r+1$
23:    end while
24:    $ans\leftarrow Agent(Q,Interaction)$triangleright Moderator agent makes the final decision
25:
26:else
27:    $ICT\leftarrow\textsc{Recruit}(Q,Complexity)$triangleright Recruit Integrated Care Team
28:    $Reports\leftarrow[]$
29:    for $Team\in ICT$  do
30:         $Report\leftarrow\textsc{GenerateReport}(Q,Team)$triangleright Each Team curates a report
31:         $Reports\leftarrow Reports+[Report]$
32:    end for
33:    $ans\leftarrow Agent(Q,Reports)$triangleright Final decision made
34:end if
35:
36:return $ans$
3.1 Query Complexity Assessment (Line 1 of Algorithm 1)
@@
To operationalize our adaptive medical decision-making framework, we first determine the complexity of the medical query using moderator LLM which functions as a generalist practitioner. The moderator LLM aims to play the role of a classifier to return the complexity level of the given medical query. The moderator LLM gets the information on how medical complexity should be defined and is instructed to classify one of three different complexity levels:
@@
    low - This level is characterized by straightforward, well-defined medical issues that can be resolved by a single primary care physician (PCP) LLM. These typically include common, acute illnesses or stable chronic conditions where the medical needs are predictable and require minimal interdisciplinary coordination.
@@
    moderate - At this level, the medical issues involve multiple interacting factors, necessitating a collaborative approach among a multidisciplinary team (MDT). These scenarios typically require the integration of diverse medical knowledge areas and coordination between specialists to develop effective care strategies.
@@
    high - This category includes highly complex medical scenarios that demand extensive coordination and combined expertise from an Integrated Cure Team (ICT). These cases often involve multiple chronic conditions, significant dependencies on medical technology, and complex decision-making that integrates various healthcare services, potentially including patient support considerations.
The method for determining complexity levels by LLM is analogous to the techniques used in real-world clinical decision-making [77, 3, 15, 63, 2]. Classification to these levels reflects the multifaceted nature of medical care, accommodating the spectrum of scenarios that healthcare professionals encounter, from routine visits to complex diagnoses that require in-depth analysis and risk assessment.
@@ 3.2 Expert Recruitment (Line 3, 7, 17 of Algorithm 1)
Given a medical query, the goal of the recruiter LLM is to recruit domain experts as individuals, in groups, or as multiple teams, based on the complexity levels determined by the moderator LLM. Specifically, we assign medical expertise and roles to multiple LLMs, instructing them to either act independently as solo medical agents or collaborate with other medical experts in a team. In Figure 8 in Appendix, we also provide frequently recruited agents for each benchmark as a reference.
@@ 3.3 Medical Collaboration and Refinement
The initial assessment protocol of our decision-making framework categorizes query complexity into low, moderate, and high. This categorization is grounded in established medical constructs such as acuity [18] for straightforward cases, comorbidity [62] and case management complexity [10] for intermediate and multi-disciplinary care requirements, and severity of illness [11] for high complexity cases requiring comprehensive management. The following outlines the specific refinement approach for each category:
low - straightforward cases (Line 2-4 of Algorithm 1)
@@
For queries classified under Low complexity, characterized by straightforward clinical decision pathways, a single primary care physician (PCP) agent (Figure 2 (a)) is deployed by the definition in Section 3.1. The domain expert who is recruited by recuriter LLM, applies few-shot prompting to the problem. The output answer, denoted as $ans$, is directly obtained from the agent’s response to $Q$ without the need for iterative refinement, formalized as $ans=Agent(Q)$, with $Agent$ representing the engaged PCP agent.
moderate - intermediate complexity cases (Line 6-14 of Algorithm 1)
@@
In addressing more complex queries, the utilization of a Multi-Disciplinary Team (MDT) (Figure 2 (b) and (c)) approach has been increasingly recognized for its effectiveness in producing comprehensive and nuanced solutions [36]. The MDT framework leverages the collective expertise of professionals from diverse disciplines, facilitating a holistic examination of the query at hand. This collaborative method is particularly advantageous in scenarios where the complexity of a problem transcends the scope of a single domain, necessitating a fusion of insights from various specialties [63, 3]. The MDT approach not only enhances decision-making quality through the integration of multidimensional perspectives but also significantly improves the adaptability and efficiency of the problem-solving process [15].
@@
Building upon this foundation, our framework specifically addresses queries of moderate complexity through a structured, multi-tiered collaborative approach. A multidisciplinary team (MDT) recruited by recruiter LLM starts an iterative discussion process aimed at reaching a consensus with at most $R$ rounds (Line 10-12). For every round, consensus within the MDT is manually verified. In the event of a disagreement, the moderator agent, consistent with the one described in Section 3.1 reviews the MDT’s discourse and formulates feedback for each agent.
high - complex care cases (Line 17-24 of Algorithm 1)
@@
In contrast to the MDT approach, the Integrated Care Team (ICT) (Figure 2 (d)) paradigm is essential for addressing the highest tier of query complexity in healthcare. This structured progression through the ICT ensures a depth of analysis that is specialized and focused at each stage of the decision-making process. Beginning with the Initial Assessment Team, moving through various diagnostic teams, and culminating with the Final Review & Decision Team, our ICT model aligns specialist insights into a cohesive narrative that informs the ultimate decision. This phased approach is implemented in Line 19-21. This phased approach, supported by evidence from recent healthcare studies, has been shown to enhance the precision of clinical decision-making, as each team builds upon the foundation laid by the previous, ensuring a meticulous and refined examination of complex medical cases [28]. The resultant reports are not only reflective of comprehensive medical evaluations but also of a systematic and layered analysis that is critical in the management of intricate health scenarios [13].
3.4 Decision Making

### v2

# MDAgents: An Adaptive Collaboration of LLMs for Medical Decision-Making

@@ Abstract
Foundation models are becoming valuable tools in medicine. Yet despite their promise, the best way to leverage Large Language Models (LLMs) in complex medical tasks remains an open question. We introduce a novel multi-agent framework, named Medical Decision-making Agents (MDAgents) that helps to address this gap by automatically assigning a collaboration structure to a team of LLMs. The assigned solo or group collaboration structure is tailored to the medical task at hand, a simple emulation inspired by the way real-world medical decision-making processes are adapted to tasks of different complexities. We evaluate our framework and baseline methods using state-of-the-art LLMs across a suite of real-world medical knowledge and medical diagnosis benchmarks. MDAgents achieved the best performance in seven out of ten benchmarks on tasks requiring an understanding of medical knowledge and multi-modal reasoning, showing a significant improvement of up to 6.5% ($p$ < 0.05) compared to previous methods’ best performances. Ablation studies reveal that MDAgents effectively determines medical complexity to optimize for efficiency and accuracy across diverse medical tasks. Notably, the combination of moderator review and external medical knowledge in group collaboration resulted in an average accuracy improvement of 11.8%. Our code can be found at https://github.com/mitmedialab/MDAgents.
@@ 1 Introduction
Medical Decision-Making (MDM) is a multifaceted and intricate process in which clinicians collaboratively navigate diverse sources of information to reach a precise and specific conclusion [92]. For instance, a primary care physician (PCP) may refer a patient to a specialist when faced with a complex case, or a patient visiting the emergency department or urgent care might be triaged and then directed to a specialist for further evaluation [51, 5]. MDM involves interpreting complex and multi-modal data, such as imaging, electronic health records (EHR), physiological signals, and genetic information, while rapidly integrating new medical research into clinical practice [74, 64]. Recently, Large Language Models (LLMs) have shown potential for AI support in MDM [68, 22, 60, 36, 46, 86]. It is known that they are able to process and synthesize large volumes of medical literature [70] and clinical information [1], as well as support probabilistic [89] and causal [38] reasoning, makes LLMs promising tools. However, there is no silver bullet in medical applications that require careful design.
Figure 1: Medical Decision-Making Agents (MDAgents) framework. Given a medical query from different medical datasets, the framework performs 1) medical complexity check, 2) recruitment, 3) analysis and synthesis, and 4) decision-making steps.
@@
While decision-making tools including multi-agent LLMs [82, 11] have shown promise in non-medical domains [31, 42, 44, 61, 32, 58], their evaluation in health applications has been limited. To date, their “generalist” design has not effectively integrated the real-world systematic MDM process [53] which requires an adaptive, collaborative, and tiered approach. Clinicians consider the current and past history of the patient, available evidence from medical literature, and their domain expertise and experience [20] for MDM. One example of MDM is to triage patients in emergency room based on the severity and complexity of their medical conditions [83, 12, 26]. Patients with pathognomonic, single uncomplicated acute conditions, or stable chronic conditions that their PCP could manage [81] could be low complexity cases. On the other hand, patients with injuries that involve multiple organs, chronic conditions with side effects, or superimposed diseases who often require multiple collaborative discussions (MDT) or sequential consultations (ICT) among specialty physicians [57, 27] are considered high complexity cases  [^1] Detailed examples of low-, moderate- and high-complexity cases are provided in Appendix E.1..
Inspired by the way that clinicians make decisions in practice, we propose Medical Decision-making Agents (MDAgents), an adaptive medical decision-making framework that leverages LLMs to emulate the hierarchical diagnosis procedures ranging from individual clinicians to collaborative clinician teams (Figure 1). MDAgents work in three steps: 1) Medical complexity check; 2) Recruitment based on medical complexity; 3) Analysis and synthesis and 4) Final decision-making to return the answer. Our contributions are threefold:
  1. 1.
    We introduce MDAgents, the first adaptive decision-making framework for LLMs that mirrors real-world MDM processes via dynamic collaboration among AI agents based on task complexity.
  2. 2.
    MDAgents demonstrate superior performance in accuracy over previous solo and group methods on 7 out of 10 medical benchmarks, and we show an effective trade-off between performance and efficiency (i.e. the number of API calls) by varying the number of agents.
  3. 3.
    We provide rigorous testing under various hyperparameters (e.g. temperatures), demonstrating better robustness of MDAgents compared to solo and group methods. Furthermore, our ablations evidence MDAgents’ ability to find the appropriate complexity level for each MDM instance.
2 Related Works
Language Models in Medical Decision-Making
LLMs have shown promise in a range of applications within the medical field [39, 59, 71, 91, 72, 14, 36, 86, 46]. They can answer questions from medical exams [41, 49], perform biomedical research [35], clinical risk prediction [36], and clinical diagnosis [63, 52]. Medical LLMs are also evaluated on generative tasks, including creating medical reports [75], describing medical images [77], constructing differentials [50], performing diagnostic dialogue with patients [73], and generating psychiatric evaluations of interviews [24]. To advance the capabilities of medical LLMs, two main approaches have been explored: (1) training with domain-specific data [28], and (2) applying inference-time techniques such as prompt engineering [63] and Retrieval Augmented Generation (RAG) [87]. While initial research has been concentrated on pre-training and fine-tuning with medical knowledge, the rise of large general-purpose LLMs has enabled training-free methods where models leverage their latent medical knowledge. For example, GPT-4 [56], with richer prompt crafting, surpasses the passing score on USMLE by over 20 points and with prompt tuning can outperform fine-tuned models including Med-PaLM [55, 54]. The promise of general-purpose models has thus inspired various techniques such as Medprompt and ensemble refinement to improve LLM reasoning [63], as well as RAG tools that use external resources to improve the factuality and completeness of LLM responses [87, 37]. Our approach leverages these techniques and the capabilities of general-purpose models while acknowledging that a solitary LLM [36, 46, 86] may not fully encapsulate the collaborative and multidisciplinary nature of real-world MDM. We thus emphasize joining multiple expert LLMs for effective collaboration in order to solve complicated medical tasks with greater accuracy.
Multi-Agent Collaboration
An array of studies have explored effective collaboration frameworks between multiple LLM agents [45, 82] to enhance capability above and beyond an individual LLM [76]. A common framework is role-playing, where each agent adopts a specific role (e.g. an Assistant Agent or a Manager Agent), a task is then broken down into sub-steps and solved collaboratively [45, 82]. While role-playing focuses on collaboration and multi-step problem-solving [84], another framework, “multi-agent debate”, prompts each agent to solve the task independently [17]. Then, they reason through other agents’ answers to converge on a shared response, this approach can improve the factuality, mathematical ability and reasoning capabilities of the multi-agent solution [17, 47]. Similar frameworks include voting [78], multi-disciplinary collaboration [68], group discussions (ReConcile [10]), and negotiating [23]. Table 1 compares existing setups across key dimensions in multi-agent interaction. Although these frameworks have shown improvement in the respective tasks, they rely on a pre-determined number of agents and interaction settings. When applied on a wider variety of tasks , this static architecture may lead to suboptimal multi-agent configurations, negatively impacting performance [48]. Furthermore, multi-agent approaches run the risk of being computationally inefficient or expensive to employ and need to justify these costs with noticable performance gains [17]. Given that different models and frameworks could generalize better to different tasks [88], we propose a framework that dynamically assigns the optimal collaboration strategy at inference time based on the complexity of the query. We apply our strategy to MDM, a task that requires teamwork and should benefit from multi-agent collaboration [68].
@@ 2 Related Works
Table 1: Comparison between our framework and previous methods (Solo and Group). Among these works, MDAgents is the only one to perform all key dimensions of LLM decision-making.
@@
Method MDAgents (Ours) Single Voting [78] Debate [17] MedAgents [68] ReConcile [10]
@@
3 MDAgents: Medical Decision-making Agents
The design of MDAgents (Figures 1 and 2) incorporates four stages: 1) Medical Complexity Check - The system evaluates the medical query, categorizing it as low, moderate, or high complexity based on clinical decision-making techniques [80, 7, 21, 67, 6]. 2) Expert Recruitment - Based on complexity, the framework activates a single Primary Care Clinician (PCC) for low complexity issues, or a Multi-disciplinary Team (MDT) or Integrated Care Team (ICT) for moderate or high complexities [43, 67, 7, 21, 33, 18]. 3) Analysis and Synthesis - Solo queries use prompting techniques like Chain-of-Thought (CoT) and Self-Consistency (SC). MDTs involve multiple LLM agents forming a consensus, while ICTs synthesize information for the most complex cases. 4) Decision-making - The final stage synthesizes all inputs to provide a well-informed answer to the medical query.
3.1 Agent Roles
Moderator.
The moderator agent functions as a general practitioner (GP) or emergency department doctor who first triages the medical query. This agent assesses the complexity of the problem and determines whether it should be handled by a single agent, a MDT, or an ICT. The moderator ensures the appropriate pathway be selected based on the query’s complexity and oversees the entire decision-making process.
Recruiter.
The recruiter agent is responsible for assembling the appropriate team of specialist agents based on the complexity assessment of the moderator. The recruiter may assign a single PCP agent for low-complexity cases, while MDT or ICT with relevant expertise will be formed for moderate and high-complexity cases.
General Doctor/Specialist.
These agents are domain-specific or general physicians recruited by the recruiter agent. Depending on the complexity of the case, they may work independently or as part of a team. General physicians handle less complex, routine cases, whereas specialists are recruited for their specific expertise in more complex scenarios. These agents engage in the collaborative decision-making process, contributing their specialized knowledge to reach a consensus or provide detailed reports for high-complexity cases.
3.2 Medical Complexity Classification (Line 1 of Algorithm 1, Appendix C.2)
@@ 3.1 Agent Roles
The first step in the MDAgents framework is to determine the complexity of a given medical query $q$ by the moderator LLM which functions as a generalist practitioner (GP). The moderator aims to act as a classifier to return the complexity level of the given medical query, it is provided with the information on how medical complexity should be defined and is instructed to classify the query into one of three different complexity levels:
@@
    Low - Simple, well-defined medical issues that can be resolved by a single PCP agent. These typically include common, acute illnesses or stable chronic conditions where the medical needs are predictable and require minimal interdisciplinary coordination.
@@
    Moderate - The medical issues involve multiple interacting factors, necessitating a collaborative approach among an MDT. These scenarios require integration of diverse medical knowledge areas and coordination between specialists through consultation to develop effective care strategies.
@@
    High - Highly complex medical scenarios that demand extensive coordination and combined expertise from an ICT. These cases often involve multiple chronic conditions, complicated surgical or trauma cases, and complex decision-making that integrates various healthcare departments and multiple specialists.
Figure 2: Illustrative example of MDAgents in a moderate complexity case from the PMC-VQA dataset. More detailed case studies can be found in Figure 11 and 12 in the Appendix.
@@ 3.3 Expert Recruitment (Line 3, 7, 17 of Algorithm 1)
Given a medical query, the goal of the recruiter LLM is to enlist domain experts as individuals, in groups, or as multiple teams, based on the complexity levels determined by the moderator LLM. Specifically, we assign medical expertise and roles to multiple LLMs, instructing them to either act independently as solo medical agents or collaborate with other medical experts in a team. In Figure 9 in the Appendix, we also provide frequently recruited agents for each benchmark as a reference.
@@ 3.4 Medical Collaboration and Refinement
The initial assessment protocol of our decision-making framework categorizes query complexity into low, moderate, and high. This categorization is grounded in established medical constructs such as acuity [25] for straightforward cases, comorbidity [65] and case management complexity [13] for intermediate and multi-disciplinary care requirements, and severity of illness [16] for high complexity cases requiring comprehensive management. We outlines the specific refinement approach:
Low - Straightforward cases (Line 2-4 of Algorithm 1).
@@
For queries classified under Low complexity, characterized by straightforward clinical decision pathways, a single PCP agent (Figure 10 (a)) is deployed by the definition in Section 3.2. The domain expert who is recruited by recuriter LLM, applies few-shot prompting to the problem. The output answer, denoted as $ans$, is directly obtained from the agent’s response to $Q$ without the need for iterative refinement, formalized as $ans=Agent(Q)$, with $Agent$ representing the engaged PCP agent.
Moderate - Intermediate complexity cases (Line 6-14 of Algorithm 1).
@@
In addressing more complex queries, the utilization of an MDT (Figure 10 (b) and (c)) approach has been increasingly recognized for its effectiveness in producing comprehensive and nuanced solutions [43]. The MDT framework leverages the collective expertise of professionals from diverse disciplines, facilitating a holistic examination of the query at hand. This collaborative method is particularly advantageous in scenarios where the complexity of a problem transcends the scope of a single domain, necessitating a fusion of insights from various specialties [67, 7]. The MDT approach not only enhances decision-making quality through the integration of multidimensional perspectives but also significantly improves the adaptability and efficiency of the problem-solving process [21].
@@
Building upon this foundation, our framework specifically addresses queries of moderate complexity through a structured, multi-tiered collaborative approach. An MDT recruited by recruiter LLM (see Figure 10 in Appendix) starts an iterative discussion process aimed at reaching a consensus with at most $R$ rounds (Line 10-12). For each round $r\in R$, agents $A_{i},i\in{1,\ldots,N}$ indicate participation and preferred interlocutors. The system facilitates message exchanges for $T$ turns. If consensus is not reached and agents agree to continue, a new round begins with access to previous conversations. For every round, consensus within the MDT is determined by parsing and comparing their opinions. In the event of a disagreement, the moderator agent, consistent with the one described in Section 3.2 reviews the MDT’s discourse and formulates feedback for each agent.
High - Complex care cases (Line 17-24 of Algorithm 1).
@@
In contrast to the MDT approach, the ICT (Figure 10 (d)) paradigm is essential for addressing the highest tier of query complexity in healthcare. This structured progression through the ICT ensures a depth of analysis that is specialized and focused at each stage of the decision-making process. Beginning with the Initial Assessment Team, moving through various diagnostic teams, and culminating with the Final Review & Decision Team, our ICT model aligns specialist insights into a cohesive narrative that informs the ultimate decision (Appendix Algorithms 1 Lines 19-21). A key component of this process is the report generation process described in Appendix with the prompt, where each team, led by a lead clinician, collaboratively produces a comprehensive report synthesizing their findings. This phased approach, supported by evidence from recent healthcare studies, has been shown to enhance the precision of clinical decision-making, as each team builds upon the foundation laid by the previous, ensuring a meticulous and refined examination of complex medical cases [33]. The resultant reports, accumulating throughout the ICT process, are not only reflective of comprehensive medical evaluations but also of a systematic and layered analysis that is critical in the management of intricate health scenarios [18].
3.5 Decision-making